What percent of 35 is 21?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(57328)

What percent of 35 is 21?
x/100 = 21/35
x = 100(21/35)
x = 60%
======================
Cheers,
Stan H.

Answer by richwmiller(9135)

35x=21
x=21/35
x=3/5
x=.60
x=60%
check
.6*35=21
3/5*35=21
3*7=21
ok